OpenStudy (anonymous):

Using the following data table and graph, calculate the average k from the data. Amount CO2 (mL) Amount of White Solid (g) y x 10 0.15 20 0.30 30 0.46 40 0.62 50 0.77 k=60, k=65, k=70, or k=64.5

OpenStudy (anonymous):

@Kash_TheSmartGuy

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

last one was y = kx. assuming this is part of the same problem, k = y/x Divide the y's by the x's to find the k's For example the first would be k = 10/0.15 = 66.67

OpenStudy (anonymous):

Then once you have all the k's, find the average by adding them up and dividing by 4

OpenStudy (anonymous):

sorry divide by 5. I miscounted the amount of k's you have to find

OpenStudy (anonymous):

I got 64.5 @peachpi

OpenStudy (anonymous):

I guess go with it. I keep getting 65.6 :/